# How Private is DP-SGD?

### v1

@@ 4.2 $\textsc{ABLQ}_{\mathcal{D}}$ vs $\textsc{ABLQ}_{\mathcal{P}}$
For $P=\mathcal{N}(\mu,\sigma^{2}I)$ and the set $E:=\{w\in\mathbb{R}^{d}:a^{\top}w-b\geq 0\}$, $P(E)=\Phi\left(\frac{a^{\top}\mu-b}{\sigma\|a\|_{2}}\right)\,.$
@@
Note that the inequality above is strict. Since $D_{e^{\varepsilon}}(P\|Q)$ is continuous in $\varepsilon$ (see Proposition C.2), there exists some $\varepsilon_{0}>0$, such that for all $\varepsilon\in[0,\varepsilon_{0})$, $\delta_{\mathcal{D}}(\varepsilon)>\delta_{\mathcal{P}}(\varepsilon)$.
@@
$E:=\textstyle\left\{w\in\mathbb{R}^{T}\big{|}\sum_{i}w_{i}>(\varepsilon+\log 2+T\log T)\sigma^{2}+{\frac{T}{2}}\right\}$
@@
$\textstyle~{}=~{}\Phi\left(-\frac{\varepsilon\sigma}{\sqrt{T}}-\frac{(T\log T+\log 2)\sigma}{\sqrt{T}}-\frac{\sqrt{T}}{2\sigma}\right)$ (3)
where we use Proposition 4.3 in the last two steps. On the other hand, we have from Proposition 3.1 that
@@
$\textstyle~{}\leq~{}\Phi\left(-\varepsilon\sigma+\frac{1}{2\sigma}\right)$ (4)
@@
by noting that for large $\varepsilon$ the most significant term inside $\Phi(\cdot)$ in Equation 5 is $-\varepsilon\sigma/\sqrt{T}$, whereas in Equation 6 the most significant term inside $\Phi(\cdot)$ is $-\varepsilon\sigma$, which decreases much faster as $\varepsilon\to\infty$, for a fixed $T$ and $\sigma$. ∎
Conjecture 4.4.
Theorem 4.2 holds for $\varepsilon_{0}=\varepsilon_{1}$.
We do not rely on this conjecture being true in the rest of this paper. We provide a numerical example that validates Theorem 4.2 and provides evidence for 4.4. In Figure 2, for $\sigma=0.3$ and $T=10$, we plot the numerically computed $\delta_{\mathcal{D}}(\varepsilon)$ (using Proposition 3.1), as well as lower and upper bounds on $\delta_{\mathcal{P}}(\varepsilon)$, computed using the open source dp_accounting library (Google’s DP Library., 2020).
4.3 $\textsc{ABLQ}_{\mathcal{P}}$ vs $\textsc{ABLQ}_{\mathcal{S}}$
$D(E_{C})$ $\textstyle~{}=~{}1-D(\mathbb{R}^{T}\smallsetminus E_{C})$
$\textstyle~{}=~{}1-\prod_{t=1}^{T}\Phi\left(\frac{C-\mu_{t}}{\sigma}\right)$
$\textstyle~{}=~{}1-\Phi\left(\frac{C-2}{\sigma}\right)\cdot\Phi\left(\frac{C}{\sigma}\right)^{T-1}$
Thus, we use the following lower bound:
In Figure 3, we set $\sigma=0.4$ and number of steps $T=10,000$ and plot $\delta_{\mathcal{D}}(\varepsilon)$, an upper bound on $\delta_{\mathcal{P}}(\varepsilon)$ (obtained using dp_accounting) and a lower bound on $\delta_{\mathcal{S}}(\varepsilon)$ as obtained via Equation 7. We find that while $\delta_{\mathcal{P}}(4)\leq 1.18\cdot 10^{-5}$, $\delta_{\mathcal{S}}(4)\geq 0.226$, that is close to $\delta_{\mathcal{D}}(4)\approx 0.244$. Even $\delta_{\mathcal{S}}(12)\geq 7.5\cdot 10^{-5}$ is larger than $\delta_{\mathcal{P}}(4)$. While the $(4,1.2\cdot 10^{-5})$-$\mathsf{DP}$ guarantee of $\textsc{ABLQ}_{\mathcal{P}}$ could have been considered as sufficiently private, $\textsc{ABLQ}_{\mathcal{S}}$ only satisfies much worse privacy guarantees. We provide additional examples in Appendix A.
4.3.1 Intuition for 3.2
  • $\bm{x}=(x_{1}=-L,\ldots,x_{n-1}=-L,x_{n}=1)$, and
  • $\bm{x}^{\prime}=(x_{1}=-L,\ldots,x_{n-1}=-L,x_{n}=\bot)$.
The case of $L>1$ is not valid, since in this case $|\psi_{t}(x)|=L>1$. However, we can still ask how well the privacy of the $n$-th example is preserved by $\textsc{ABLQ}_{\mathcal{B}}$, by considering the hockey stick divergence between $\textsc{ABLQ}_{\mathcal{B}}(\bm{x})$ and $\textsc{ABLQ}_{\mathcal{B}}(\bm{x}^{\prime})$.
In summary, the main differing aspect about $\textsc{ABLQ}_{\mathcal{S}}$ and $\textsc{ABLQ}_{\mathcal{P}}$ is that in the former, the non-differing examples can leak information about the location of the differing example in the shuffled order, whereas that is not the case in the latter. While we sketched an argument that works asymptotically as $L\to\infty$, we see glimpses of it already at $L=1$.
Our intuitive reasoning behind 3.2 is that even in the case of (vector valued) adaptive query methods, in order to “leak the most information” about the differing record $x_{i}$ between $\bm{x}$ and $\bm{x}^{\prime}$, it seems natural that the query $\psi_{t}(x_{j})$ should evaluate to $-\psi_{t}(x_{i})$ for all $j\neq i$. If the query method satisfies this condition for all $t$, then it is easy to show that $(P_{\mathcal{S}},Q_{\mathcal{S}})$ tightly dominates $(\textsc{ABLQ}_{\mathcal{S}}(\bm{x}),\textsc{ABLQ}_{\mathcal{S}}(\bm{x}^{\prime}))$. 3.2 then asserts that this is indeed the worst case.
5 Conclusion & Future Directions
There are several interesting directions that remain to be investigated. In our work, we provide a technique to only provide a lower bound on the privacy guarantee when using a shuffle batch sampler. It would be interesting to have a tight accounting method for $\textsc{ABLQ}_{\mathcal{S}}$. A first step towards this could be to establish 3.2, which if true, might make numerical accountants for computing the hockey stick divergence possible. While this involves computing a high dimensional integral, it might be possible to approximate using importance sampling; e.g., such approaches have been attempted for $\textsc{ABLQ}_{\mathcal{P}}$ (Wang et al., 2023).
However, it remains to be seen how the utility of DP-SGD would be affected when we use the correct privacy analysis for $\textsc{ABLQ}_{\mathcal{S}}$ instead of the analysis for $\textsc{ABLQ}_{\mathcal{P}}$, which has been used extensively so far and treated as a good “approximation”. Alternative approaches such as $\mathsf{DP}\text{-}\mathsf{FTRL}$ (Kairouz et al., 2021; McMahan et al., 2022) that do not rely on amplification might turn out to be better if we instead use the correct analysis for $\textsc{ABLQ}_{\mathcal{S}}$, in the regimes where such methods are currently reported to under perform compared to DP-SGD.
Another important point to note is that the model of shuffle batch sampler we consider here is a simple one. There are various types of data loaders used in practice, which are not necessarily captured by our model. For example tf.data.Dataset.shuffle takes in a parameter of buffer size $b$. It returns a random record among the first $b$ records, and immediately replaces it with the next record ($(b+1)$-th in this case), and continues repeating this process. This leads to an asymmetric form of shuffling. Such batch samplers might thus require more careful privacy analysis.
The notion of DP aims to guarantee privacy even in the worst case. For example in the context of DP-SGD, it aims to protect privacy of a single record even when the model trainer and all other records participating in the training are colluding to leak information about this one record. And moreover, releasing the final model is assumed to leak as much information as releasing all intermediate iterates. Such strong adversarial setting might make obtaining good utility to be difficult. Alternative techniques for privacy amplification such as amplification by iteration (Feldman et al., 2018; Altschuler & Talwar, 2022) or through convergence of Langevin dynamics (Chourasia et al., 2021) have been studied, where only the last iterate of the model is released. However, these analyses rely on additional assumptions such as convexity and smoothness of the loss functions. Investigating whether it is possible to relax these assumptions to make amplification by iteration applicable to non-convex models is an interesting future direction.
Appendix A Additional Evaluations
A.1 $\varepsilon$ vs $\sigma$ for fixed $\delta$ and $T$
We first plot $\varepsilon$ vs $\sigma$ for fixed $\delta$ and $T$. We compute upper bounds on $\varepsilon_{\mathcal{P}}(\delta)$ using Rényi differential privacy (RDP) as well as using privacy loss distributions (PLD). These accounting methods are provided in the open source Google dp_accounting library (Google’s DP Library., 2020).
  • In Figure 4, we set $\delta=10^{-6}$ and number of steps $T=10,000$. In particular, for $\sigma=0.5$, $\varepsilon_{\mathcal{P}}(\delta)\leq 1.96$ (as per PLD accounting) and $\varepsilon_{\mathcal{P}}(\delta)\leq 3.43$ (as per RDP accounting), whereas on the other hand, $\varepsilon_{\mathcal{S}}(\delta)\geq 10.994$ and $\varepsilon_{\mathcal{D}}(\delta)\approx 10.997$.
  • In Figure 5, we set $\delta=10^{-5}$ and number of steps $T=1000$. In particular, for $\sigma=0.7$, $\varepsilon_{\mathcal{P}}(\delta)\leq 0.61$ (as per PLD accounting) and $\varepsilon_{\mathcal{P}}(\delta)\leq 1.64$ (as per RDP accounting), whereas on the other hand, $\varepsilon_{\mathcal{S}}(\delta)\geq 6.528$ and $\varepsilon_{\mathcal{D}}(\delta)\approx 6.652$.
A.2 $\delta$ vs $\varepsilon$ for fixed $\sigma$ and $T$.
Next, we plot $\delta$ vs $\varepsilon$ for fixed $\sigma$ and $T$. We compute upper bounds on $\delta_{\mathcal{P}}(\varepsilon)$ using Rényi differential privacy (RDP) as well as using privacy loss distributions (PLD).
Appendix B Proof of Theorem 4.1
We make a simple observation that shuffling the dataset first cannot degrade the privacy guarantee of any mechanism.
Lemma B.1.
Proof.
If $\mathcal{M}$ satisfies $(\varepsilon,\delta)$-$\mathsf{DP}$, we have that for any measurable event $E\subseteq\mathcal{O}$, and any $\pi$, it holds that $P_{\pi}(E)-e^{\varepsilon}Q_{\pi}(E)\leq\delta$, since $\bm{x}_{\pi}$ and $\bm{x}^{\prime}_{\pi}$ are also adjacent datasets. Hence we have that
$D_{e^{\varepsilon}}(P\|Q)$ $~{}=~{}\sup_{E}P(E)-e^{\varepsilon}Q(E)~{}=~{}\sup_{E}\sum_{\pi}\frac{1}{n!}\left(P_{\pi}(E)-e^{\varepsilon}Q_{\pi}(E)\right)~{}\leq~{}\sup_{E}\max_{\pi}P_{\pi}(E)-e^{\varepsilon}Q_{\pi}(E)~{}\leq~{}\delta$
Thus, $\mathcal{M}_{\mathcal{S}}$ also satisfies $(\varepsilon,\delta)$-DP. ∎
Proof of Theorem 4.1.
Appendix C Additional proofs for $\textsc{ABLQ}_{\mathcal{D}}$ vs $\textsc{ABLQ}_{\mathcal{P}}$
Proposition C.1.
For all $T\in\mathbb{N}$ and distributions $A,B$, $D_{1}(A^{\otimes T}\|B^{\otimes T})\leq 1-(1-D_{1}(A\|B))^{T}$. And hence $D_{1}(A^{\otimes T}\|B^{\otimes T})\leq T\cdot D_{1}(A\|B)$ and equality can occur only if $T=1$ or $D_{1}(A\|B)=0$.
Proof.
$\Pr[(X_{1},\dots,X_{T})\neq(Y_{1},\dots,Y_{T})]$
$=1-\Pr[(X_{1},\dots,X_{T})=(Y_{1},\dots,Y_{T})]$
$=1-\Pr[X=Y]^{T}.$
Proposition C.2.
For $\varepsilon_{1}<\varepsilon_{2}$, it holds that $D_{e^{\varepsilon_{1}}}(P\|Q)-D_{e^{\varepsilon_{2}}}(P\|Q)\leq e^{\varepsilon_{2}}-e^{\varepsilon_{1}}$
Proof.
We have that
$~{}=~{}e^{\varepsilon_{2}}-e^{\varepsilon_{1}}\,.$

### v2

@@ Appendix A Additional Evaluations
For $P=\mathcal{N}(\mu,\sigma^{2}I)$ and the set $E:=\{w\in\mathbb{R}^{d}:a^{\top}w-b\geq 0\}$, it holds that $P(E)=\Phi\left(\frac{a^{\top}\mu-b}{\sigma\|a\|_{2}}\right)\,.$
Proposition C.2.
For all $T\in\mathbb{N}$ and distributions $A,B$, it holds that $D_{1}(A^{\otimes T}\|B^{\otimes T})\leq 1-(1-D_{1}(A\|B))^{T}$. And hence $D_{1}(A^{\otimes T}\|B^{\otimes T})\leq T\cdot D_{1}(A\|B)$ and equality can occur only if $T=1$ or $D_{1}(A\|B)=0$.
Proof.
$\Pr[(X_{1},\dots,X_{T})\neq(Y_{1},\dots,Y_{T})]$
$=1-\Pr[(X_{1},\dots,X_{T})=(Y_{1},\dots,Y_{T})]$
$=1-\Pr[X=Y]^{T}.$
Proposition C.3.
For $\varepsilon_{1}<\varepsilon_{2}$, it holds that $D_{e^{\varepsilon_{1}}}(P\|Q)-D_{e^{\varepsilon_{2}}}(P\|Q)\leq e^{\varepsilon_{2}}-e^{\varepsilon_{1}}$.
Proof.
We have that
$~{}=~{}e^{\varepsilon_{2}}-e^{\varepsilon_{1}}\,.\qed$
@@
Note that the inequality is strict for $T>1$. Since $D_{e^{\varepsilon}}(P\|Q)$ is continuous in $\varepsilon$ (see Proposition C.3), there exists some $\varepsilon_{0}>0$, such that for all $\varepsilon\in[0,\varepsilon_{0})$, $\delta_{\mathcal{D}}(\varepsilon)>\delta_{\mathcal{P}}(\varepsilon)$.
@@
$E:=\textstyle\left\{w\in\mathbb{R}^{T}\big{|}\sum_{i}w_{i}>(\varepsilon+\log 2+T\log T)\sigma^{2}+{\frac{T}{2}}\right\}.$
@@
$\textstyle~{}=~{}\Phi\left(-\frac{\varepsilon\sigma}{\sqrt{T}}-\frac{(T\log T+\log 2)\sigma}{\sqrt{T}}-\frac{\sqrt{T}}{2\sigma}\right),$ (4)
where we use Proposition C.1 in the last two steps. On the other hand, we have from Proposition 3.1 that
@@
$\textstyle~{}\leq~{}\Phi\left(-\varepsilon\sigma+\frac{1}{2\sigma}\right).$ (5)
@@
by noting that for large $\varepsilon$ the most significant term inside $\Phi(\cdot)$ in (6) is $-\varepsilon\sigma/\sqrt{T}$, whereas in (7) the most significant term inside $\Phi(\cdot)$ is $-\varepsilon\sigma$, which decreases much faster as $\varepsilon\to\infty$, for a fixed $T>1$ and $\sigma>0$. ∎